# Acceleration Radiation of Freely Falling Atoms in Bardeen Regular Black Hole Spacetimes

### v1

@@ I Introduction
The deep interplay between gravitation, quantum field theory, and thermodynamics was first revealed in the seminal discovery that black holes radiate thermally with a temperature proportional to their surface gravity GibbonsHawking; Hawking. In Hawking’s calculation, quantum fields on a fixed black hole background produce a flux of particles at infinity, while in the Unruh effect, uniformly accelerated observers in flat spacetime detect a thermal bath in the Minkowski vacuum Unruh:1976db; FullingUnruh. These phenomena are now understood as cornerstone examples of quantum field theory in curved spacetime and non-inertial frames BirrellDavies; WaldQFT; WaldGR; DeWittQFT.
@@
A complementary and operational perspective on these effects is provided by particle-detector models, in which localized two-level systems (or Unruh-DeWitt detectors) couple to quantum fields along prescribed worldlines. Such detectors have been used to probe Unruh and Hawking thermality in a variety of settings, including uniformly accelerated trajectories, cosmological spacetimes, and black hole geometries, with an emphasis on response functions, entanglement harvesting, and the weak equivalence principle Tjoa:2018xre; Lopp:2018lxl; You:2018qgr; Fulling:2018lez; Chakraborty:2019ltu; MasoodASBukhari:2023flr.
@@
Within this framework, a particularly fruitful line of research is the quantum-optics program initiated by Scully and collaborators, who recast Hawking-Unruh physics in terms of atoms interacting with quantum fields in curved backgrounds and accelerated frames Scully:2017utk; Svidzinsky:2018jkp; Svidzinsky:2019jqr. A central result of this program is the notion of horizon-brightened acceleration radiation (HBAR): when a two-level atom falls freely into a Schwarzschild black hole through a Boulware-like vacuum in the presence of a near-horizon mirror, it can still become excited while emitting radiation with a Planckian spectrum characterized by the Hawking temperature, despite the absence of outgoing Hawking flux at future null infinity Scully:2017utk; Camblong:2020pme. This effect arises from the relative acceleration between the atom and the field modes constrained by the mirror, and can be traced to a universal near-horizon conformal quantum mechanics (CQM) structure in the radial Klein-Gordon equation Camblong:2020pme; Azizi:2020gff; Azizi:2021qcu; Azizi:2021yto; Das:2022qcr; Das:2022qpx; Good:2021ffo; Ordonez:2025sqp.
@@
The near-horizon CQM perspective shows that, for a wide class of static and stationary black holes, the dominant physics of s-wave modes is governed by an inverse-square potential whose effective coupling is fixed by the surface gravity. This insight has led to a series of extensions of the HBAR setup: to rotating and charged black holes Azizi:2020gff; Sen:2023zfq, to general classes of static spherically symmetric metrics beyond Schwarzschild Sen:2022cdx, to braneworld black holes Das:2023rwg, and to backgrounds with external fields, such as dark energy or dark matter distributions Bukhari:2022wyx; Bukhari:2023yuy. More recent work has pushed the formalism to derivative-coupled detectors and modified gravity scenarios Pantig:2025okn; Das:2025rzz; Rahaman:2025grm; Rahaman:2025mrr; Tang:2025eew, to Lorentz-violating spacetimes, and resonantly driven setups where redshift-enhanced and Floquet-type acceleration radiation can emerge Pantig:2025igg; Masood:2024glj; Wang:2025jvb; Pan:2024jix.
@@
A parallel development has focused on the thermodynamic and entropic aspects of acceleration radiation. Building on master-equation techniques and the CQM structure, Azizi et al. have shown how the coarse-grained dynamics of a single cavity mode interacting with a flux of infalling atoms leads to entropy production and area-law-like relations that mirror Hawking’s original thermodynamic picture Azizi:2021qcu; Azizi:2021yto. Subsequent works have explored HBAR entropy and its quantum-gravity-motivated corrections in quantum-corrected and renormalization-group improved black hole backgrounds Sen:2022tru; Jana:2024fhx; Jana:2025hfl; Ovgun:2025isv; Eissa:2025vhs, including inverse-logarithmic and generalized-uncertainty-principle (GUP) corrections, as well as causal-diamond geometries. These studies strengthen the view that the HBAR setup provides a versatile quantum-optical laboratory for probing horizon thermality, the equivalence principle, and possible quantum-gravity imprints Chatterjee:2021fue; Chatterjee:2021fsg; Barman:2021oum; Barman:2021kwg; Liu:2019soq; Barman:2023rhd; Allen:2019muo.
@@
At the same time, there is growing interest in connecting these theoretical developments to experiments and analogue systems. Proposals range from Ramsey interferometry and cavity-QED schemes designed to witness acceleration radiation Costa:2020asn; Lopes:2021ajs; Gregori:2023tun, to studies of dynamical Casimir and gravitational-wave-induced effects in quantum optical settings Sorge:2018zfd; Sorge:2023tyn; Chatterjee:2022pkj, to analogue-gravity platforms based on water waves and optical media that mimic horizon kinematics and surface gravity BarceloVisserLiberati; Rozenman:2024ymh; Xu:2021cbw. Together with more formal advances on nonlocal fields, nonvacuum states, and Wigner distributions in Rindler spacetime Prokhorov:2019sss; Das:2022qcr; Das:2022qpx; Dubey:2024abz; Stargen:2025siq, these efforts suggest that acceleration radiation and HBAR may provide experimentally relevant windows into quantum aspects of spacetime.
@@
However, most HBAR studies to date have focused on singular black holes or on effective quantum-corrected metrics that still retain a central curvature singularity, albeit softened or surrounded by a new structure Camblong:2020pme; Sen:2022tru; Sen:2022cdx; Jana:2024fhx; Ovgun:2025isv. In contrast, regular black holes–and in particular the Bardeen solution, which arises as an effective geometry of Einstein gravity coupled to nonlinear electrodynamics—provide de Sitter-like cores with no curvature singularity at the origin Bambi:2013ufa; Ayon-Beato:2000mjt. The Bardeen parameter $g$ controls the size of the regular core and interpolates between a Schwarzschild-like regime ($g\to 0$) and an extremal zero-temperature remnant in which the outer and inner horizons coincide. Regular black holes are widely used as phenomenological models of short-distance quantum-gravity effects and as candidate endpoints of Hawking evaporation. However, their imprint on acceleration radiation, detector response, and HBAR entropy has not been systematically explored.
@@
In this work, we bring the HBAR and near-horizon CQM program to the realm of Bardeen regular black holes. We consider a neutral scalar field in the Bardeen geometry and a two-level atom freely falling through a Boulware-like vacuum in the presence of a stretched-horizon mirror, following the quantum-optics formulation of Refs. Scully:2017utk; Camblong:2020pme; Azizi:2021qcu; Azizi:2021yto; Ordonez:2025sqp. Our goals are twofold. First, we show that the near-horizon CQM structure responsible for HBAR thermality persists in the Bardeen spacetime, with an effective inverse-square potential whose coupling is fixed by the Bardeen surface gravity and hence by the core parameter $g$. This leads to a Planckian excitation spectrum for the falling atom characterized by the Bardeen Hawking temperature and a strong suppression of the radiation as the extremal, cold-remnant limit is approached. Second, we construct a coarse-grained master equation and an associated HBAR entropy for a selected cavity mode in this regular background, and investigate how singularity resolution and the de Sitter-like core deform the Wien-type relations and area-law behavior found in earlier HBAR entropy analyses Azizi:2021qcu; Azizi:2021yto; Sen:2022tru; Jana:2024fhx; Jana:2025hfl; Ovgun:2025isv; Eissa:2025vhs. In this way, the Bardeen parameter acts as a tunable “regularization scale” controlling both the strength of acceleration radiation and the structure of its entropy, providing a quantum-optical probe of regular black hole remnants and their thermodynamic properties.
@@ II Bardeen black hole spacetime
We consider a static, spherically symmetric, regular (nonsingular) black hole of Bardeen type. Its line element can be written in standard Schwarzschild-like coordinates as Bambi:2013ufa; Ayon-Beato:2000mjt
@@ III Geodesic equations
Near the event horizon $\mathcal{H}$, with $r\sim r_{g}$, it is convenient to introduce the shifted coordinate $x=r-r_{g}$, which measures the radial distance to the horizon. The near-horizon expansions of $F(r)$ and its derivatives can then be organized as a hierarchical series in powers of $x$ Tang:2025eew:
@@ IV Near-horizon conformal quantum mechanics equation
We now turn to the quantum field that interacts with the falling atom Azizi:2021qcu; Azizi:2021yto. We consider a minimally coupled, massless scalar field obeying the Klein-Gordon equation
@@
We expand the field in modes as Camblong:2020pme
@@
Near the horizon $\mathcal{H}$, using the expansions (13), the radial equation simplifies drastically. The leading contributions in the potential behave as an inverse square in the proper distance to the horizon, and the radial equation takes the CQM form Camblong:2003mz
@@ V Particle spectrum of acceleration radiation
Since no Hawking flux reaches the asymptotic observer due to the mirror, the field is taken to be in a Boulware-like vacuum Boulware:1974dm,
@@
where $\tau$ is the proper time along the atomic trajectory and $|1_{\nu}\rangle$ denotes the one-photon mode with frequency $\nu$. In natural units, the interaction Hamiltonian in the interaction picture can be written as ScullyHBAR
@@
For $s\gg\sigma$ (e.g. in the geometrical optics regime $\omega\gg\nu$ so that $s\simeq\omega/e$), we can extend the upper limit to infinity and evaluate the integral by analytic continuation of the standard identity Camblong:2020pme
@@ VI HBAR entropy and area law in the Bardeen geometry
In this section we adapt the quantum-optics derivation of HBAR entropy à la Scully et al. and the conformal-quantum-mechanics (CQM) framework of Refs. Scully:2017utk; Azizi:2021qcu; Azizi:2021yto; Ordonez:2025sqp to the Bardeen geometry. The result is an HBAR entropy flux that satisfies the same area law as in the Schwarzschild case, but with all geometric dependence encoded in the Bardeen surface gravity and horizon radius. We emphasize the role of the Bardeen parameter $g$ and show that both the HBAR energy and entropy fluxes are quenched in the extremal Bardeen limit.
@@ VI.3 Wien’s displacement law in the Bardeen background
The Planckian spectrum (47) for the HBAR excitation probability implies the existence of a Wien-type displacement law in the Bardeen geometry. As in ordinary blackbody radiation, there is a characteristic wavelength at which the spectrum attains its maximum, and the product of this wavelength with the effective temperature is a universal constant determined solely by the functional form of the spectrum ScullyHBAR; Das:2023rwg; Jana:2024fhx. In the present context, the relevant temperature is the Bardeen Hawking temperature $T_{H}^{(\mathrm{B})}(g)$, which encodes all dependence on the regular core parameter $g$ through the surface gravity. In this subsection we make this statement precise by recasting the excitation probability in wavelength space, deriving the corresponding displacement law, and analyzing how the spectral peak responds to variations of $g$.

### v2

# Acceleration Radiation of Freely Falling Atoms: Nonlinear Electrodynamic Effects

@@ I Introduction
The deep interplay between gravitation, quantum field theory, and thermodynamics was first revealed in the seminal discovery that black holes radiate thermally with a temperature proportional to their surface gravity Gibbons and Hawking (1977); Hawking (1975). In Hawking’s calculation, quantum fields on a fixed black hole background produce a flux of particles at infinity, while in the Unruh effect, uniformly accelerated observers in flat spacetime detect a thermal bath in the Minkowski vacuum Unruh (1976); Fulling (1973). These phenomena are now understood as cornerstone examples of quantum field theory in curved spacetime and non-inertial frames Birrell and Davies (1982); Wald (1995, 1984); DeWitt (1979).
@@
A complementary and operational perspective on these effects is provided by particle-detector models, in which localized two-level systems (or Unruh-DeWitt detectors) couple to quantum fields along prescribed worldlines. Such detectors have been used to probe Unruh and Hawking thermality in a variety of settings, including uniformly accelerated trajectories, cosmological spacetimes, and black hole geometries, with an emphasis on response functions, entanglement harvesting, and the weak equivalence principle Tjoa et al. (2018); Lopp et al. (2018); You and Li (2018); Fulling and Wilson (2019); Chakraborty and Majhi (2019); Masood A. S. Bukhari and Wang (2024).
@@
Within this framework, a particularly fruitful line of research is the quantum-optics program initiated by Scully and collaborators, who recast Hawking-Unruh physics in terms of atoms interacting with quantum fields in curved backgrounds and accelerated frames Scully et al. (2018a); Svidzinsky et al. (2018); Svidzinsky (2019). A central result of this program is the notion of horizon-brightened acceleration radiation (HBAR): when a two-level atom falls freely into a Schwarzschild black hole through a Boulware-like vacuum in the presence of a near-horizon mirror, it can still become excited while emitting radiation with a Planckian spectrum characterized by the Hawking temperature, despite the absence of outgoing Hawking flux at future null infinity Scully et al. (2018a); Camblong et al. (2020). This effect arises from the relative acceleration between the atom and the field modes constrained by the mirror, and can be traced to a universal near-horizon conformal quantum mechanics (CQM) structure in the radial Klein-Gordon equation Camblong et al. (2020); Azizi et al. (2021a, b, c); Das and Majhi (2022); Das et al. (2023); Good and Linder (2022); Ordonez et al. (2025).
@@
The near-horizon CQM perspective shows that, for a wide class of static and stationary black holes, the dominant physics of s-wave modes is governed by an inverse-square potential whose effective coupling is fixed by the surface gravity. This insight has led to a series of extensions of the HBAR setup: to rotating and charged black holes Azizi et al. (2021a); Sen et al. (2023), to general classes of static spherically symmetric metrics beyond Schwarzschild Sen et al. (2022b), to braneworld black holes Das et al. (2024), and to backgrounds with external fields, such as dark energy or dark matter distributions Bukhari et al. (2023); Bukhari and Wang (2024). More recent work has pushed the formalism to derivative-coupled detectors and modified gravity scenarios Pantig and Övgün (2025); Das et al. (2025); Rahaman (2025b, a); Tang et al. (2025), to Lorentz-violating spacetimes, and resonantly driven setups where redshift-enhanced and Floquet-type acceleration radiation can emerge Pantig et al. (2025); Masood (2024); Wang et al. (2025); Pan and Zhang (2024).
@@
A parallel development has focused on the thermodynamic and entropic aspects of acceleration radiation. Building on master-equation techniques and the CQM structure, Azizi et al. have shown how the coarse-grained dynamics of a single cavity mode interacting with a flux of infalling atoms leads to entropy production and area-law-like relations that mirror Hawking’s original thermodynamic picture Azizi et al. (2021b, c). Subsequent works have explored HBAR entropy and its quantum-gravity-motivated corrections in quantum-corrected and renormalization-group improved black hole backgrounds Sen et al. (2022a); Jana et al. (2024, 2025); Övgün and Pantig (2026); Eissa et al. (2025), including inverse-logarithmic and generalized-uncertainty-principle (GUP) corrections, as well as causal-diamond geometries. These studies strengthen the view that the HBAR setup provides a versatile quantum-optical laboratory for probing horizon thermality, the equivalence principle, and possible quantum-gravity imprints Chatterjee et al. (2021b, a); Barman and Majhi (2021); Barman et al. (2022); Liu and Xiong (2019); Barman and Majhi (2024); Allen (2020).
@@
At the same time, there is growing interest in connecting these theoretical developments to experiments and analogue systems. Proposals range from Ramsey interferometry and cavity-QED schemes designed to witness acceleration radiation Costa et al. (2020); Lopes et al. (2021); Gregori et al. (2024), to studies of dynamical Casimir and gravitational-wave-induced effects in quantum optical settings Sorge and Wilson (2018); Sorge (2024); Chatterjee and Majumdar (2022), to analogue-gravity platforms based on water waves and optical media that mimic horizon kinematics and surface gravity Barcelo et al. (2005); Rozenman et al. (2024); Xu and Wang (2021). Together with more formal advances on nonlocal fields, nonvacuum states, and Wigner distributions in Rindler spacetime Prokhorov et al. (2020); Das and Majhi (2022); Das et al. (2023); Dubey and Kolekar (2025); Stargen and Sudhir (2025), these efforts suggest that acceleration radiation and HBAR may provide experimentally relevant windows into quantum aspects of spacetime.
@@
However, most HBAR studies to date have focused on singular black holes or on effective quantum-corrected metrics that still retain a central curvature singularity, albeit softened or surrounded by a new structure Camblong et al. (2020); Sen et al. (2022a, b); Jana et al. (2024); Övgün and Pantig (2026). Lorentzian–Euclidean black holes provide a nonsingular framework for avoiding singularities, where atemporality (rooted in conservation laws) crucially influences the causal structure, null-geodesic dynamics, and matter accretion Capozziello et al. (2024); De Bianchi et al. (2025); Capozziello et al. (2025). In contrast, regular black holes–and in particular the Bardeen solution, which arises as an effective geometry of Einstein gravity coupled to nonlinear electrodynamics—provide de Sitter-like cores with no curvature singularity at the origin Bambi and Modesto (2013); Ayon-Beato and Garcia (2000). The Bardeen parameter $g$ controls the size of the regular core and interpolates between a Schwarzschild-like regime ($g\to 0$) and an extremal zero-temperature remnant in which the outer and inner horizons coincide. Regular black holes are widely used as phenomenological models of short-distance quantum-gravity effects and as candidate endpoints of Hawking evaporation. However, their imprint on acceleration radiation, detector response, and HBAR entropy has not been systematically explored.
@@
In this work, we bring the HBAR and near-horizon CQM program to the realm of Bardeen regular black holes. We consider a neutral scalar field in the Bardeen geometry and a two-level atom freely falling through a Boulware-like vacuum in the presence of a stretched-horizon mirror, following the quantum-optics formulation of Refs. Scully et al. (2018a); Camblong et al. (2020); Azizi et al. (2021b, c); Ordonez et al. (2025). Our goals are twofold. First, we show that the near-horizon CQM structure responsible for HBAR thermality persists in the Bardeen spacetime, with an effective inverse-square potential whose coupling is fixed by the Bardeen surface gravity and hence by the core parameter $g$. This leads to a Planckian excitation spectrum for the falling atom characterized by the Bardeen Hawking temperature and a strong suppression of the radiation as the extremal, cold-remnant limit is approached. Second, we construct a coarse-grained master equation and an associated HBAR entropy for a selected cavity mode in this regular background, and investigate how singularity resolution and the de Sitter-like core deform the Wien-type relations and area-law behavior found in earlier HBAR entropy analyses Azizi et al. (2021b, c); Sen et al. (2022a); Jana et al. (2024, 2025); Övgün and Pantig (2026); Eissa et al. (2025). In this way, the Bardeen parameter acts as a tunable “regularization scale” controlling both the strength of acceleration radiation and the structure of its entropy, providing a quantum-optical probe of regular black hole remnants and their thermodynamic properties.
@@ II Bardeen black hole spacetime
We consider a static, spherically symmetric, regular (nonsingular) black hole of Bardeen type. Its line element can be written in standard Schwarzschild-like coordinates as Bambi and Modesto (2013); Ayon-Beato and Garcia (2000)
@@ III Geodesic equations
Near the event horizon $\mathcal{H}$, with $r\sim r_{g}$, it is convenient to introduce the shifted coordinate $x=r-r_{g}$, which measures the radial distance to the horizon. The near-horizon expansions of $F(r)$ and its derivatives can then be organized as a hierarchical series in powers of $x$ Tang et al. (2025):
@@ IV Near-horizon conformal quantum mechanics equation
We now turn to the quantum field that interacts with the falling atom Azizi et al. (2021b, c). We consider a minimally coupled, massless scalar field obeying the Klein-Gordon equation
@@
We expand the field in modes as Camblong et al. (2020)
@@
Near the horizon $\mathcal{H}$, using the expansions (13), the radial equation simplifies drastically. The leading contributions in the potential behave as an inverse square in the proper distance to the horizon, and the radial equation takes the CQM form Camblong and Ordonez (2003)
@@ V Particle spectrum of acceleration radiation
Since no Hawking flux reaches the asymptotic observer due to the mirror, the field is taken to be in a Boulware-like vacuum Boulware (1975),
@@
where $\tau$ is the proper time along the atomic trajectory and $|1_{\nu}\rangle$ denotes the one-photon mode with frequency $\nu$. In natural units, the interaction Hamiltonian in the interaction picture can be written as Scully et al. (2018b)
@@
For $s\gg\sigma$ (e.g. in the geometrical optics regime $\omega\gg\nu$ so that $s\simeq\omega/e$), we can extend the upper limit to infinity and evaluate the integral by analytic continuation of the standard identity Camblong et al. (2020)
@@ VI HBAR entropy and area law in the Bardeen geometry
In this section we adapt the quantum-optics derivation of HBAR entropy à la Scully et al. and the conformal-quantum-mechanics (CQM) framework of Refs. Scully et al. (2018a); Azizi et al. (2021b, c); Ordonez et al. (2025) to the Bardeen geometry. The result is an HBAR entropy flux that satisfies the same area law as in the Schwarzschild case, but with all geometric dependence encoded in the Bardeen surface gravity and horizon radius. We emphasize the role of the Bardeen parameter $g$ and show that both the HBAR energy and entropy fluxes are quenched in the extremal Bardeen limit.
@@ VI.3 Wien’s displacement law in the Bardeen background
The Planckian spectrum (47) for the HBAR excitation probability implies the existence of a Wien-type displacement law in the Bardeen geometry. As in ordinary blackbody radiation, there is a characteristic wavelength at which the spectrum attains its maximum, and the product of this wavelength with the effective temperature is a universal constant determined solely by the functional form of the spectrum Scully et al. (2018b); Das et al. (2024); Jana et al. (2024). In the present context, the relevant temperature is the Bardeen Hawking temperature $T_{H}^{(\mathrm{B})}(g)$, which encodes all dependence on the regular core parameter $g$ through the surface gravity. In this subsection we make this statement precise by recasting the excitation probability in wavelength space, deriving the corresponding displacement law, and analyzing how the spectral peak responds to variations of $g$.